OpenStudy (anonymous):

The volume of a rectangular prism is equal to x3 + 4x2 + 3x. What is the longest edge of the prism? x + 1 x + 3 4x x3

Directrix (directrix):

x3 + 4x2 + 3x = l*w*h x*(x^2 + 4x + 3) = l*w*h x* ( x + 3) * (x + 1) = l*w*h x has to be positive --> x + 3 is the longest edge